### v1

# Optimizations of electronic stopping cross section measurements using the backscattering method

Abstract
The accurate determination of electronic stopping cross sections is essential for advancing nuclear science and materials research. In this work, we present an optimized methodology for stopping power measurements using the backscattering technique, with a focus on minimizing both random and systematic uncertainties. By systematically analyzing the sources of uncertainty and refining the experimental geometry, we establish a robust framework for high-precision measurements. Our approach was applied to the stopping power of helium ions in gold thin films, demonstrating excellent agreement with reference models such as SRIM and ICRU-49. The results reveal that carefully chosen measurement angles can effectively balance statistical precision and systematic accuracy, achieving total uncertainties below 3% across a wide energy range. This study provides a refined strategy for stopping power determination, offering valuable improvements for ion beam analysis and benchmarking theoretical models.
keywords:
Stopping cross section , Backscattering method , Uncertainty optimization , Systematic uncertainty
journal: Nuclear Instruments and Methods, Section B
1 Introduction
The electronic stopping power of ions in matter plays a crucial role in nuclear science and its applications. Although theoretical understanding of the physical processes involved in energy loss phenomena is advancing [1, 2, 3], experimental works focusing on traceable experimental measurements become crucial to benchmark models. Accurate experimental data also help to improve semiempirical approaches to produce tables that are used in simulations targeting a wide range of applications.
In this work, we demonstrate the optimization of the electronic stopping power measurements to minimize the random and systematic uncertainties in the backscattering method. Our approach consists of estimating the sources of uncertainties and improving the method by choosing the best geometry that favors the reduction of uncertainties.
2 Methods
2.1 The back-scattering method
There are several approaches to measuring stopping power using backscattering data [4]. In this work, we adopt the method based on determining the peak width in thin-film measurements, as described in [5]. Fig. 1 shows a typical spectrum of a pure metallic film deposited on a light substrate. The peak is formed by all scattered ions from the surface down to the film-substrate interface. Its width ($\Delta E$) is the difference between the energy of ions scattered by the film atoms at the surface ($K\cdot E_{0}$, with $K$ being the kinematic factor and $E_{0}$ the incidence beam energy) and the exit energy of ions re-emerging after being scattered by film atoms at the interface ($E_{out}$). Assuming a homogeneous stopping in the test film along both, the entry and exit paths, the energy difference can be written as:
Figure 1: Illustrative spectrum for the RBS technique.
The peak width can be determined by adding the energy loss during the path of the ion into the film and the energy loss during its exit from a thin film of thickness $t$. Assuming homogeneous stopping power along both the entry and exit paths, this expression can be written as:
$\Delta E=\frac{K\cdot t}{\cos(\theta_{in})}[\epsilon]_{in}+\frac{t}{\cos(\theta_{out})}[\epsilon]_{out}$ (1)
with:
$\epsilon_{j}=\left.\frac{1}{N}\frac{dE}{dx}\right|_{j}$ (2)
being the stopping cross section in the $j$ path, either the entry (index $in$) or the exit (index $out$) of the film. $\theta_{in}$ and $\theta_{out}$ represent the angles between the ion’s entry and exit paths and the surface normal, respectively. $N$ is the material atomic density.
The fundamental principle of this methodology is that by measuring the peak width in two different geometrical configurations (i.e. two pairs of values of $\theta_{in}$ and $\theta_{out}$) for the same incident energy, one can obtain a system of equations. The solutions for $[\epsilon_{in}]$ and $[\epsilon_{out}]$ yield the stopping cross section at two different mean energies. In matrix formalism, this can be expressed as:
$\begin{pmatrix}\Delta E^{A}\\ \\ \Delta E^{B}\\ \end{pmatrix}=\begin{pmatrix}\frac{K\cdot t}{\cos(\theta_{in}^{A})}&\frac{t}{\cos(\theta_{out}^{A})}\\ \\ \frac{K\cdot t}{\cos(\theta_{in}^{B})}&\frac{t}{\cos(\theta_{out}^{B})}\\ \end{pmatrix}\begin{pmatrix}[\epsilon]_{in}\\ \\ [\epsilon]_{out}\ \\ \end{pmatrix}$ (3)
where index $A$ and $B$ denote the two different measurements. To better constrain the problem, we impose the conditions $\theta_{in}^{A}=\theta_{out}^{B}=\Theta_{1}$ and $\theta_{in}^{B}=\theta_{out}^{A}=\Theta_{2}$. This assumption is necessary to reduce the number of variables, thereby enhancing result visualization and simplifying the experimental procedure. It is important to note that the factor $K$ implicitly depends on the scattering angle, which remains fixed throughout the analysis. As a consequence, the experimental procedure involves tilting the sample in such a way that, in the first measurement, the incidence angle is $\Theta_{1}$ and the exit angle is $\Theta_{2}$, while in the second measurement, the angles are reversed, with the incidence angle set to $\Theta_{2}$ and the exit angle to $\Theta_{1}$.
There are different approximations to determine the mean energy at the entry and exit paths of the film [5]. Here, we adopt the following approximations:
$\begin{split}\bar{E}_{in}&=E_{0}-\frac{1}{4}\Delta\bar{E}\\ \bar{E}_{out}&=K\cdot E_{0}-\left(\frac{2\cdot K+1}{4}\right)\Delta\bar{E}\end{split}$ (4)
with $\Delta\bar{E}=(\Delta E^{A}+\Delta E^{B})/2$. This is an alternative definition based on the symmetrical mean energy approximation proposed in [5], which is more convenient for an experimental approach.
Both the assumption of homogeneous energy loss and the approximations used to calculate the mean energy along the ion’s entry and exit paths in the film introduce systematic errors. These assumptions do not account for the non-linearity of the stopping cross-section curve. The impact of these approximations on the final result will be evaluated in Section 2.3.
2.2 Accounting for random uncertainties
From Eq. 3, we can propagate the uncertainties in the determination of $\Delta E_{j}$ to the values of $[\epsilon_{j}]$ using the definition of a covariance matrix:
$V_{\tilde{A}}=\left(X^{T}V^{-1}X\right)^{-1}$ (5)
with:
$X=\begin{pmatrix}\frac{K\cdot t}{\cos(\Theta_{1})}&\frac{t}{\cos(\Theta_{2})}\\ \\ \frac{K\cdot t}{\cos(\Theta_{2})}&\frac{t}{\cos(\Theta_{1})}\\ \end{pmatrix}$ (6)
$V=\begin{pmatrix}{\sigma_{A}}^{2}&0\\ \\ 0&{\sigma_{B}}^{2}\\ \end{pmatrix}$ (7)
Let $\sigma_{A}$ and $\sigma_{B}$ be the uncertainties associated with the measurements of $\Delta E_{A}$ and $\Delta E_{B}$, respectively. In this work, we assumed $\sigma_{A}=\sigma_{B}=G/\sqrt{6}$, where $G$ is the gain calibration of the spectra (see A). This provides a conservative estimate of the precision at which the peak width can be determined from histogram data.
The adjusted values of the stopping cross-section of a pair of measurements are then given by Eq. 8.
$\begin{pmatrix}[\epsilon]_{in}\\ \\ [\epsilon]_{out}\ \\ \end{pmatrix}=\left(X^{T}V^{-1}X\right)^{-1}\cdot X^{T}\cdot V^{-1}\cdot\begin{pmatrix}\Delta E^{A}\\ \\ \Delta E^{B}\\ \end{pmatrix}$ (8)
Note that arranging the matrix $X$ in this format implies alternating $\Theta_{1}$ and $\Theta_{2}$ as angles of the entry and exit paths in two consecutive measurements at the same incidence beam energy.
Eq. 5 results in a matrix whose diagonal holds the uncertainty of our system:
$\sigma_{[\epsilon]_{in}}=\frac{G}{\sqrt{6}Kt}\cdot\frac{\sqrt{\sec^{2}(\Theta_{1})+\sec^{2}(\Theta_{2})}}{|\sec^{2}(\Theta_{1})-\sec^{2}(\Theta_{2})|}$ (9)
$\sigma_{[\epsilon]_{out}}=\frac{G}{\sqrt{6}t}\cdot\frac{\sqrt{\sec^{2}(\Theta_{1})+\sec^{2}(\Theta_{2})}}{|\sec^{2}(\Theta_{1})-\sec^{2}(\Theta_{2})|}$ (10)
and the outer diagonal results in the covariance between the $[\epsilon]_{in}$ and $[\epsilon]_{out}$ values, which is used to calculate the correlation factor r as:
$\textbf{r}=\frac{\textbf{cov}([\epsilon]_{in},[\epsilon]_{out})}{\sigma_{[\epsilon]_{in}}\cdot\sigma_{[\epsilon]_{out}}}=-{}\frac{2\sec(\Theta_{1})\sec(\Theta_{2})}{\sec^{2}(\Theta_{1})+\sec^{2}(\Theta_{2})}$ (11)
Note that the correlation factor is always negative, whatever the combination of $\Theta_{1}$ and $\Theta_{2}$.
2.3 Accounting for systematic error
To estimate the systematic uncertainties, we tested Eqs. 4 and 8 against the SRIM curve. The peak width, $\Delta E$, has been computed directly using the integral (as in Eq. 12) form and SRIM data, while the methodology described in Section 2.1 is applied to determine $[\epsilon]$ using Eq. 8. These values are then compared to the SRIM values themselves interpolated at the mean energies given by Eq. 4. The differences between the interpolated and the output of the matrix calculation are considered as systematic errors. This accounts for systematic errors inserted either by the surface approximation or by the calculation of the mean energy in the entry and exit paths.
[…]
This approach is justified because SRIM data provide a reliable approximation of the stopping values in light of the current knowledge status of this process. A new measurement aims at improving this knowledge status by restricting the uncertainties. Since our objective is to minimize uncertainties, this method helps identify geometries where our assumptions remain valid.
$s({[\epsilon]_{j}})=[\epsilon]_{j}-\textbf{SRIM}(\bar{E}_{j})$ (13)
with $s({[\epsilon]_{j}})$ being the systematic error and $\textbf{SRIM}(\bar{E_{j}})$ being the SRIM interpolation at the energy $\bar{E_{j}}$. $j$ is an index that denotes $in$ or $out$.
2.4 Finding the best geometry
With the definitions of random and systematic uncertainties for the case of determination of $[\epsilon]$, the total uncertainty can be defined as:
$\sigma_{total}=\sqrt{s({[\epsilon]_{j}})^{2}+\sigma_{[\epsilon]_{j}}^{2}}$ (14)
Note that the common factor in the statistical uncertainty of $[\epsilon]_{in}$ and $[\epsilon]_{out}$, as in Eq. 9 and 10, can be written as the function $f(\Theta_{1},\Theta_{2})$:
$f(\Theta_{1},\Theta_{2})=\frac{\sqrt{\sec^{2}(\Theta_{1})+\sec^{2}(\Theta_{2})}}{|\sec^{2}(\Theta_{1})-\sec^{2}(\Theta_{2})|}$ (15)
which diverges when $\Theta_{1}\approx\Theta_{2}$. In special, when $\Theta_{1}=\Theta_{2}$ the system in Eq. 3 becomes undetermined. $f(\Theta_{1},\Theta_{2})$ reaches minimum at $\Theta_{1}=0$ and $\Theta_{2}=\pi/2$ or $\Theta_{1}=\pi/2$ and $\Theta_{2}=0$. However, as $\Theta_{1}$ or $\Theta_{2}$ approaches $\pi/2$, the length of the corresponding ion’s path at entry or exit increases, leading to greater systematic errors.
Therefore, an optimal choice of $\Theta_{1}$ and $\Theta_{2}$ must balance reducing random uncertainties while also minimizing systematic errors. Since systematic uncertainty cannot be evaluated analytically, this procedure must be performed numerically.
3 Results and discussion
To illustrate the application of the method, we calculated both the random and systematic uncertainties, along with the correlation factor, to optimize an experiment aiming at measuring the electronic stopping cross-section of gold for helium ions. The adopted film thickness in these calculations was 45 nm. The results for each contribution to the total uncertainty are presented in the panel of Fig. 2.
Figure 2: Panel showing different contributions to the final uncertainty calculated assuming the measurement of helium stopping cross-section in a $3\times 10^{17}$ at./cm2 (45 nm) thick gold film. The incident alpha beam energy is 1000 keV and the detector was placed at 120 scattering angle. Top line shows the relative random uncertainty calculated for the way-in (left) and way-out (center) the film, and the correlation factor of these both values (right). The bottom line shows systematic uncertainties for the way-in (left) and way-out (center) the film. Color scale in uncertainty plots are in logarithm scale and saturated at 10% for improved visualization. The color scale in the correlation factor map in linear and restricted between -1 and 0 since it is always negative.
As mentioned above, the two types of uncertainty exhibit opposite behavior, and the correlation factor presents only negative values (overestimating in the stopping cross section during the ion’s entry into the film leads to underestimation during its exit, and vice versa). The correlation factor weakens if $\Theta_{1}>60^{\circ}$ and $\Theta_{2}<20^{\circ}$. Uncorrelated measurements are only achievable at specific angle combinations that lead to high systematic errors and are therefore not recommended.
The combination of random and systematic uncertainties yields the total uncertainties as shown in Fig. 3. We also calculated the final uncertainties for incident energies of 500 keV and 5000 keV, with the results presented in Figs. 4 and 5, respectively.
From these results, it is clear that, to keep the total uncertainty below 3% in the overall energy range, $\Theta_{1}$ should be kept between 50 and 80 and $\Theta_{2}$ should be below 30. Furthermore, it can be observed that the total uncertainties below 1% are not feasible in this experiment. If higher accuracy is needed, alternative methods should be explored [6].
Figure 3: Heat maps showing calculated total uncertainty assuming the measurement of helium stopping cross-section in a $3\times 10^{17}$ at./cm2 (45 nm) thick gold film. The incident energy is 1000 keV and detector placed at 120 scattering angle. Total uncertainty in the way-in (left) and way-out (right). Color scale in uncertainty plots are in logarithm scale and saturated at 10% for improved visualization.
Figure 4: Heat maps showing calculated total uncertainty assuming the measurement of helium stopping cross-section in a $3\times 10^{17}$ at./cm2 (45 nm) thick gold film. The incident energy is 500 keV and detector placed at 120 scattering angle. Total uncertainty in the way-in (left) and way-out (right). Color scale in uncertainty plots are in logarithm scale and saturated at 10% for improved visualization.
Figure 5: Heat maps showing calculated total uncertainty assuming the measurement of helium stopping cross-section in a $3\times 10^{17}$ at./cm2 (45 nm) thick gold film. The incident energy is 5000 keV and detector placed at 120 scattering angle. Total uncertainty in the way-in (left) and way-out (right). Color scale in uncertainty plots are in logarithm scale and saturated at 10% for improved visualization.
Finally, we performed experiments with incident energies ranging from 500 keV up to 5100 keV. The set of angles was chosen as $\Theta_{1}=0^{\circ}$ and $\Theta_{2}=60^{\circ}$. This choice results in a maximum total uncertainty of 3% for measurements with incident energy of 500 keV and a minimum of 2% for measurements with incident energy of 1000 keV (at the maximum of the stopping cross-section curve). The correlation factor in this set of angles is approximately 0.7.
Fig. 6 shows the experimental data obtained using our method compared to SRIM and ICRU-49 data. The black points in the plot represent additional experimental data from the IAEA database. In the upper plot, we observe the overall agreement of our data with the semi-empirical models, while in the bottom plot, we display the relative residuals with SRIM as the reference. We provide densely spaced experimental data to better define the statistical limits of the method. At lower energies, it is evident that all experimental data fall within 4% in residuals amplitude, a value that is very close to the final uncertainty calculated using the method proposed here.
Figure 6: Stopping cross section measured for gold to helium ions (upper plot) and Relative residuals calculated using SRIM as reference (bottom plot). The geometry used was defined using the methodology reported here. 1 TFU = $1\times 10^{15}$ at./cm2. In the residuals, data was labeled according to the way used in the average, i.e. way in or way out the thin film.
In the residuals plot in Fig. 6, our experimental data were labeled accordingly to determine whether they were obtained using the entry path (green) or the exit path (red). We observe a great consistency of the data determined by both paths.
The comparison with the models shows reasonable agreement in the limit of uncertainty bars both for SRIM and ICRU-49, with the latter better capturing the fine structure of data distribution in the residuals plot.
4 Conclusions
In this work, we have presented a novel methodology for optimization when measuring stopping power using backscattering data, offering an accurate and reliable approach for determining the stopping cross-section in thin films. Our method, which is based on a systematic treatment of uncertainties, provides a robust framework for minimizing random and systematic errors while optimizing geometries for precise measurements.
The experimental results, which were obtained for helium in gold in a wide range of incident energies, show good agreement with well-established semi-empirical models, such as SRIM and ICRU-49, and highlight the effectiveness of our approach in achieving uncertainties that are close to the theoretical limits. The careful selection of geometries and the use of densely spaced data points were crucial in refining the statistical accuracy of our measurements, particularly at lower energies.
Although total uncertainties below 1% were not feasible in this experiment, the method provides a reliable tool to plan stopping cross-section measurements, allowing better restriction of the spread in obtained experimental values.
Acknowledgments
The authors thank the financial support provided by CNPq-INCT-FNA (project number 464898/2014-5). TFS acknowledges the Brazilian funding agency CNPq (project number 406982/2021-0). FM acknowledges CNEN (Project No. 2020.06.IPEN.32). This research used resources from the Laboratory for Materials Analysis with Ion Beams - LAMFI-USP, of the University of São Paulo. The authors acknowledge laboratory staff for their assistance during the experiments.
During the preparation of this work, the authors used ChatGPT by OpenAI to improve language and readability. After using this tool, the authors reviewed and edited the content as needed and assume full responsibility for the content of the publication.
Appendix A Random uncertainties in $\Delta E$ determination
To determine the thin film peak width at half of its maximum, it is necessary to determine in which channels the half height occurs in the trailing and falling edges. Being conservative, we can assume a uniform probability distribution within one channel, leading to $\sigma_{\texttt{edge}}=1/\sqrt{12}$. With $\Delta E$ calculated as the difference in channels of the trailing and falling edges times the gain factor $G$, the uncertainty of $\Delta E$ can be calculated as:
$\sigma_{\Delta E}=G\cdot\sqrt{\left(\frac{1}{\sqrt{12}}\right)^{2}+\left(\frac{1}{\sqrt{12}}\right)^{2}}=\frac{G}{\sqrt{6}}$ (16)